Graph Type: Hyperbola
Explanation: Graph comes close to x=0 and y=-0 (asymptotes) without touching the lines

What happens to y as x gets larger?
What happens to x as y gets larger?
d) Given a line containing the points (1,4), (2,7), and (3,10) determine the slope-intercept form of the equation, provide one additional point on this line, and graph the function.
Equation in Slope-Intercept Form:
Show your work here:
Give one additional point in (x,y) form that would fall on this line:


Answer by stanbon(75887)

Hyperbola
Explanation: Graph comes close to x=0 and y=-0 (asymptotes) without touching the lines
For example, the equation could be y=1/x
-------------
What happens to y as x gets larger?
y gets smaller.
-----------
What happens to x as y gets larger?
x gets smaller
----------------
d) Given a line containing the points (1,4), (2,7), and (3,10) determine the slope-intercept form of the equation, provide one additional point on this line, and graph the function.
Form: y=mx+b; need to find m and b.
m=[7-4]/[2-1]=3
4=3(1)+b
b=1
Equation in Slope-Intercept Form:
y=3x+1
Give one additional point in (x,y) form that would fall on this line:
(0,1)
Cheers,
stan H.